OpenStudy (anonymous):

Find the slope of a line perpendicular to 3x+5y=15 5/3 -3/5 -5/3 3/5 * I got -3/5

Nnesha (nnesha):

what is slope of your line that they gave you

Nnesha (nnesha):

and remember perpendicular should be opposite and change sign for example if you r slope is a/b then your perpendicular slope would be -b/a flip it and change sign

OpenStudy (nurali):

3X + 5Y = 15 5Y = -3X + 15 Y = (-3/5)X + 3 The slope of this line is -3/5. The slope of a line perpendicular to it is 5/3.

OpenStudy (compassionate):

Hey. So basically, a perpendicular slope will have an oposite slope. Our first step, of course, would be solving the equation into y = mx + b; so simply transpose your terms to solve for y. 3x+5y=15 i. Subtract 3x (3x - 3x + 5y = 15 - 3x) ii. Now you have 5y = 15 - 3x Okay, so from here just solve for y. As for the slope; your slope will always be opposite. Everything stays the same; just take the opposite of your slope. So if I have 2/1; the perpendicular form will be 1/2

OpenStudy (compassionate):

 Notice how the slopes are opposite of one another!

OpenStudy (anonymous):

ahh ok

OpenStudy (compassionate):

Parallel lines have equal slopes Perpendicular lines have opposites